Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-hours and its standard cost card per unit is as follows:

Direct material: 5 pounds at $8.00 per pound $ 40.00
Direct labor: 2 hours at $14 per hour 28.00
Variable overhead: 2 hours at $5 per hour 10.00
Total standard variable cost per unit $ 78.00

The company also established the following cost formulas for its selling expenses:

Fixed Cost per Month Variable Cost per Unit Sold
Advertising $ 200,000
Sales salaries and commissions $ 100,000 $ 12.00
Shipping expenses $ 3.00

The planning budget for March was based on producing and selling 25,000 units. However, during March the company actually produced and sold 30,000 units and incurred the following costs:

  1. Purchased 160,000 pounds of raw materials at a cost of $7.50 per pound. All of this material was used in production.
  2. Direct-laborers worked 55,000 hours at a rate of $15.00 per hour.

  3. Total variable manufacturing overhead for the month was $280,500.

  4. Total advertising, sales salaries and commissions, and shipping expenses were $210,000, $455,000, and $115,000, respectively.

12. What amounts of advertising, sales salaries and commissions, and shipping expenses would be included in the company’s flexible budget for March?
Answer #1

12) Amount in Flexible budget

Advertising = 200000

Sales salaries and commission = (30000*12+100000) = $460000

Shipping expense = 30000*3 = 90000